OpenStudy (anonymous):

Graph the systems of linear inequalities. 1. y > -6x – 2 y > -10 + 5x y≤0

OpenStudy (anonymous):

Do you need too shade?

OpenStudy (anonymous):

no just graph

OpenStudy (anonymous):

Oh okay then well what we know is that the equations are in this form y=mx+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So lets see the first one. y > -6x – 2 y=mx+b, remember that m=slope and that b is the y-intercept. Plot the y-intercept first.

OpenStudy (anonymous):

I dont know how to use the graph on here

OpenStudy (anonymous):

 Use this

OpenStudy (anonymous):

sorry for the poor art but to use it hit draw

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then copy the previous drawing and draw the line on it okay?

OpenStudy (anonymous):